3. Gains from trade Consider two neighboring island countries called Felicidad and Bellissima. They each have 4 labor hours available per week that they can use to produce corn, jeans, or a combination of both. The following table shows the amount of corn or jeans that can be produced using 1 hour of labor. Corn Country Felicidad Bellissima Jeans (Pairs per hour of labor) 20 (Bushels per hour of labor) Initially, suppose Bellissima uses 1 million hours of labor per week to produce corn and 3 million hours per week to produce jeans, while Felicidad uses 3 million hours of labor per week to produce corn and 1 million hours per week to produce jeans. Consequently, Felicidad produces 15 million bushels of corn and 20 million pairs of jeans, and Bellissima produces 8 million bushels of corn and 48 million pairs of jeans. Assume there are no other countries willing to trade goods, so, in the absence of trade between these two countries, each country consumes the amount of corn and jeans it produces. 1/4. 1/2. 2, or 4 pair Felicidad opportunity cost of producing 1 bushel of corn is 1/4. 1/2, 2, or 4 pair of jeans, and Bellissimas opportunity cost of producing 1 bushel of corn is bellissima, felicidad bellissima, felicidad of jeans. Therefore has a comparative advantage in the production of corn, ana has a comparative advantage in the production of jeans. Suppose that each country completely specializes in the production of the good in which it has a comparative advantage, producing only that good. In this case, the country that produces corn will produce million bushels per week, and the country that produces jeans will produce million pairs per week. In the following table, select the amount of each good that each country exports and imports in the boxes across the row marked Trade Action, and enter each countrys final consumption of each good on the line marked Consumption. When the two countries did not specialize, the total production of corn was 23 million bushels per week, and the total production of jeans was 68 million pairs per week. Because of specialization, the total production of corn has increased by production of jeans has increased by million bushels per week, and the total million pairs per week. Because the two countries produce more corn and more jeans under specialization, each country is able to gain from trade. Calculate the gains from trade-that is, the amount by which each country has increased its consumption of each good relative to the first row of the table. In the following table, enter this difference in the boxes across the last row (marked Increase in Consumption) Felicidad Bellissima Corn Jeans Corn eans (Millions of bushels) (Millions of pairs) (Millions of bushels) (Millions of pairs) Without Trade Production 15 20 48 Consumption 15 20 48 With Trade Production Trade action Consumption export 18-import 18 ▼ export 54, import 54 ▼ export 18, import 18 ▼ export 54, import 54 ▼ Gains from Trade Increase in Consumption
Answer #1

Felicidad's opportunity cost of producing 1 bushel of corn is 20/5 = 4 pairs of jeans and Bellissima's opportunity cost of producing 1 bushel of corn is 16/8 =2 pairs of jeans. Because Bellissima has a lower opportunity cost for the production of corn. Therefore, Bellissima has a comparative advantage in the production of corn and Felicidad has a comparative advantage in the production of jeans.

Suppose that each country completely specializes in the production of the good in which it has a comparative advantage , producing only that good . In this case,the country that produces corn i.e Bellissima will produce (4)(8) = 32 million bushels per month And the country that produces jeans i.e Felicidad will produce 4(20)=80 million pairs per month.

Suppose the country that produces corn trades 18 million bushels of corn to the other country in exchange for 54 million pairs of jeans.

Felicidad Bellissima
Corn (milions of bushels) Jeans (Millions of pairs) Corn (Millions of bushels) Jeans (Millions of pairs)
Without trade Production 15 20 8 48
Consumption 15 20 8 48
With Trade Production 0 80 32 0
Trade action Import 18 Export 54 Export 18 Import 54
Consumption 18 (80-54)=26 (32-18)=14 54
Gains from trade Increase in consumption (18-15)= 3 (26-20)= 6 (14-8)=6 (54-48)=6

When the two countries did not specialize ,the total production of corn was 23 million bushels per month and the total production of jeans was 68 millions pairs per month. Because of specialization ,the total production of corn was 32 million bushels per month and total production of jeans was 80 million pairs per month. This implies that because of specialization , total production of corn has increased by (32-23)=9 million bushels per month and total production of jeans has increased by (80-68)=12 million pairs per month.